OpenStudy (anonymous):

solve for system of equation 3x-7y=5 & 9y=5x+5 a (-3,2) b (8,5) c (-10,-5) d (0,0)

OpenStudy (sleepyjess):

c (-10, -5) For this you can just plug in the numbers since it is multiple choice. The first number is always going to be x and the second y. 3(-10)-7(-5)=5 -30+35=5 5=5 9(-5)=5(-10)+5 -45=-50+5 -45=-45